# Density-Based Long-Range Electrostatic Descriptors for Machine Learning Force Fields

### v1

@@ III.1 Long-Range Effects of Point Charges
As an initial test set for the present long-range model, a gas of randomly distributed point charges, analogous to the one used in Grisafi and Ceriotti (2019), was constructed. The set comprises systems of varying volumes, each containing 64 atoms, with 32 having a positive charge of $+1$ and the other 32 having a negative charge of $-1$. To prevent large energies the minimum distance between two atoms was set to $2.5$. The training and validation data were generated using the Ewald energies and forces from VASP. Only long-range descriptors were used for this system.
@@
Figure 1: The RMSPE in the forces for various values of the minimal $\zeta_{\text{min}}=\zeta_{\text{max}}/s^{N_{\text{max}}-1}$ of the radial basis functions, scaling constants $s$, and number of radial basis functions $N_{\text{max}}$ for the gas of point charges. It is evident that the error reduction is primarily associated with a decreasing scaling constant. All curves with a small scaling constant of $s=2$ yield a minimum at $\zeta_{\text{min}}\approx 0.09$. Exponents larger than the minimum value are insufficiently long-ranged to accurately capture long-range electrostatic interactions.
@@
Overall, reducing the scaling constant $s$ reduces the RMSPE. This is because a smaller scaling constant results in a denser distribution of radial basis functions, allowing for better a description of the electrostatic kernel ($1/r$). For the electrostatic interactions to be accurately described, it is necessary to have a sufficient number of basis functions such that they decay sufficiently slowly. We see a more or less pronounced minimum around $\zeta_{\text{min}}\approx 0.09$. The second minimum to the left corresponds to the case that the second smallest $\zeta$ becomes approximately $0.09$. For $\zeta$, the inclusion of values smaller than $0.09$ is unnecessary but also does not significantly degrade the quality of the fit.
The plot shows a steep increase in error on the right-hand side, indicating that the smallest exponent is insufficient to capture long-range electrostatic interactions. This occurs when the radial basis function with the smallest exponent decays too quickly and the basis functions are unable to model the long-range interactions.
@@
The values $N_{\text{max}}=6$, $s=2$ and $\zeta_{\text{max}}=2$ are the optimal choices among all possible values for the density-based descriptors for the gas of point charges. These hyperparameters will be used in the subsequent calculations. Similar tests must be conducted to determine the optimal set of these three hyperparameters for other materials. However, the values obtained here provide a good starting point, which can help avoid extensive grid searches.
Next, we compare our approach with the LODE method, which is implemented in the way presented in II.4. It is known from theory that the two-body LODE descriptor is ideal for a system where only Coulomb interactions are present Grisafi and Ceriotti (2019). Consequently, using only two-body descriptors and a single radial basis function for a small cutoff is sufficient. In this case one descriptor for each interaction pair (+1, +1), (+1, -1), (-1, +1), and (-1, -1) suffices, as these exactly represent the interaction of one positive or negative point charge with all other positive or negative point charges. In our case, the radial basis function, onto which the charge is projected, is a spherical Bessel function that is smoothly cut off at the cutoff radius.
@@
Figure 2 shows the learning curves for the two different methods. The LODE method was tested with various combinations of cutoff radii $R_{\text{cut}}$ and the number of radial basis functions $N_{\text{max}}$, resulting in different numbers of expansion coefficients, and potentially probing the electrostatic potentials at points further away from the central atom. For the LODE model, this adds irrelevant information for the simple point charge model.
@@
where $\textbf{r}_{i}$ and $\textbf{r}_{j}$ are the positions of the particles and $Z_{i}$ and $Z_{j}$ are their charges. Since the descriptors include information about each atom type $J$ independently the charges are not relevant for the implementation and can be set to one. Then the regression will essentially determine $Z_{i}Z_{j}$.
@@
is obtained as a function of the electrostatic potential $\textbf{r}_{i}$
@@
This potential is exactly "measured" by the LODE descriptor with one radial basis function if the cutoff radius is smaller than the minimal atomic distance (the Gaussian broadening adds additional width to the source charges tough). The use of a cutoff sphere is possible due to the Gauss law. According to this, the electrostatic potential at the center of the sphere can be determined by a probe charge of finite size, as long as there is no source term inside the sphere.
@@
With a cutoff radius of 3 Å, the value measured by the single descriptor does not correspond exactly to the value of the potential at the center of the sphere. In this case, the electrostatic potential of point charges that enter the projection sphere can not determined accurately. This is why the blue line levels off at relatively high RMSPEs. Generally, considering the Gaussian broadening $\sigma=0.3$ of the atomic source charges, a cutoff radius of $R_{\text{cut}}\lesssim 2.5~{}$Å$-5\sigma$ must be chosen, where 2.52.52.5~{}2.5Å is the shortest nearest neighbor distance in our models and a factor 5 before $\sigma$ assures that the broadened charge has decayed to negligible values.
@@
When using four radial basis functions (turquoise line), a linear combination of the values of the descriptors can approximate a $\delta$-function at the origin, although the model requires additional training data to learn the precise linear combination that corresponds best to the $\delta$ function.
@@
When training only on energies and predicting forces (as shown in Fig. 2(c) and (d)), the energy predictions are mostly more accurate. However, the errors in the predicted forces are notably higher and seem to stagnate at some point. The same precision as in the mixed training on energies and forces is never reached. This indicates that the models require information about the forces to make highly accurate predictions.
@@
A comparison of our density-based descriptor with the ideal LODE descriptor for this material reveals that the LODE descriptor is superior. While our density-based descriptor yields satisfactory results, it is necessary to employ more radial basis functions (specifically, six instead of one) and perform a hyperparameter search to identify the optimal choice of $\zeta_{\text{max}}$, scaling constant $s$, and the number of radial basis functions. This demonstrates that our density-based descriptor, as predicted by theory, is not the best possible descriptor for this prototypical system. Nevertheless, it is still capable of accurately describing the interactions, albeit requiring more training data.
@@
It is noteworthy that the LODE descriptor with one radial basis function and a cutoff radius smaller than the minimal atomic distance (red line) and the density-based descriptor (black line) can achieve errors below 1% when sufficient training data are used. The LODE descriptor requires only about 40 training structures to achieve a relative error of 1%. The density-based descriptor requires 250 training structures to achieve the same accuracy and then stagnates. We believe the stagnation at very small errors is likely related to the condition number of the design problem becoming very large, which makes it difficult for the pseudo-inversion to separate "noise" from relevant data. Nevertheless, both models give very satisfactory results. These are better than those typically obtained using machine learning, as the errors for MLFFs are typically in the mid-single digit percentage range (around 2 to 10%). The hyperparameters, such as the number of radial basis functions and reciprocal lattice points, can be adjusted to increase the speed and memory efficiency of our computations while still achieving excellent results.
III.2 Sodium Chloride
@@
After demonstrating the models’ ability to describe systems with purely long-range interactions, they were applied to real materials. Liquid sodium chloride (NaCl) was chosen as the first material due to the significant difference in electronegativity between the Na and Cl atoms, which suggests the presence of non-negligible long-range electrostatic interactions. The combined descriptors capture both short- and long-range interactions, enabling analysis of material properties depending on both types of interactions.
@@
The dataset for this material consists of 1014 different structures, each containing 64 Na and 64 Cl atoms. The configurations are taken from VASP molecular dynamics (MD) simulations with 50000 steps and a time step of 1.5 fs. During this run the material was heated from 1100K to 1400K using the Langevin thermostat. An energy cutoff of 350 eV was used. The PAW potentials used were "PAW_PBE Na_pv 19Sep2006" for Na and "PAW_PBE Cl 06Sep2006" for Cl, considering seven valence electrons for each element.
@@
Table 1 presents the results for NaCl obtained using five different methods: purely local/short-range descriptors (Local), MACE without MP (MACE no MP), LODE combined with short-range descriptors (LODE), long-range density-based descriptors combined with short-range descriptors (Density) and MACE with one MP layer (MACE). The combination of short- and long-range descriptors follows the procedure outlined in Eq. (15).
@@
For the short-range part of the descriptor, the hyperparameters were set to $\sigma=0.5$, $R_{\text{cut}}=~{}$6 Å, $N_{\text{max}}=6$ and $L_{\text{max}}=3$ where $L_{\text{max}}$ corresponds to the angular quantum number and indicates the number of angular basis functions. For the density-based part, the optimal choice was $N_{\text{max}}=7$, $s=2$, $\zeta_{\text{max}}=1.5$ and $\sigma=0.3$. For the LODE part $R_{\text{cut}}=$ 1 Å, $N_{\text{max}}=1$ and $\sigma=0.3$ were used. A polynomial kernel of order $\zeta=4$ was used.
@@
For MACE, only invariant features, $R_{\text{cut}}=~{}$6 Å, descriptors with up to four-body interactions, force weights of $1000$, energy weights of $10$ and $100$ epochs were used. After $80$ epochs the energy weights are increased to lower the error in the energies.
@@
Table 1: The root-mean-square percentage error (RMSPE) was calculated for the validation data of liquid NaCl using combined density and LODE descriptors as well as local-only descriptors and MACE Batatia et al. (2022) with and without Message Passing (MP). Radial cutoffs of typically 6 Å are used. The results show that our density-based approach outperforms the LODE method for real materials and is as effective as MACE with MP. Nevertheless, including long-range interactions is essential as the purely local scheme and MACE without MP yield large errors.
Local MACE no MP LODE Density MACE
RMSPE E (%) 7.7 8.5 8.1 2.2 3.0
RMSPE F (%) 12.6 8.5 8.7 3.3 3.1
The density-based descriptors and Message Passing MACE outperform the purely local method. LODE performs similarly to MACE with no MP.
When comparing the LODE method and the density-based approach, it is evident that the density-based descriptors perform better. This is because LODE is optimal for purely electrostatic interactions, yet lacks the capability to describe intermediate range interactions. In real materials, the $Z_{i}Z_{j}/\|\textbf{r}_{ij}\|=Z_{i}Z_{j}/r_{ij}$ coulomb interaction between particles $i$ and $j$ is screened by the electrons, resulting in an effective interaction
@@
At very large distances the interaction will be screened by the ion-damped macroscopic dielectric constant $\epsilon^{-1}$. However, at medium distance, the electronic screening is less pronounced. LODE is designed to model $1/r$ interactions but not screened interactions $V$. The present approach is far more flexible.
@@
Even when the cutoff radius $R_{\text{cut}}$ of the purely local method is systematically increased to make the descriptors longer-ranged the result never becomes quite as good as for the combined density method or MACE. This is shown in Fig. 3. These data were calculated using the MLFF implemented in VASP. For the three-body descriptors a cutoff radius of $5$ was used since this yields better results than 6 Å. This is also the reason why the initial value in this plot is better than the one shown in Tab. 1 obtained using polipy4vasp. In polipy4vasp a distinction between two-body and three-body cutoffs is not possible and the best compromise is found to be 6 Å. The initial error (9.8%) is comparable to MACE without MP (8%) but requires us to perform a hyperparameter search for the cutoff of the three-body terms. As shown in Fig. 3, the error decreases as the cutoff of the two-body descriptors increases, reaching a value slightly below 5 % when the cutoff is sufficiently large. In summary, naively increasing the two-body cutoff does not quite allow us to recover the accuracy of the long-range models. Nevertheless, the present observations have prompted us to routinely use relatively larger cutoffs for the two-body descriptors (8 Å), and fairly small radial cutoffs for the three-body descriptors (5 Å) in VASP.
@@
Figure 3: RMSPE in the forces as a function of the two-body cutoff radius $R_{\text{cut}}$ for NaCl for the short-range descriptors. Data was calculated using VASP. Even when the cut-off radius is increased to 12 Å  the local method is unable to match the accuracy of the combined density method or MACE.
As with the hyperparameter optimization demonstrated for the gas of point charges, the hyperparameters for NaCl also required optimization. During this process, we observed that the inclusion of unnecessary information, such as additional long-range radial basis functions with exponents that decay more slowly than necessary for the description of electrostatic interactions in this material, negatively impacts the predictive capability of the model. The addition of radial basis functions that do not encode relevant physical information to the descriptor results in a distortion of the similarity measure of the kernel. Descriptors that have a high degree of similarity with respect to the kernel measure may exhibit a reduction in this similarity when the additional, irrelevant information is added. Thus, the inclusion of irrelevant information reduces the learning efficiency.
@@
Furthermore, the inclusion of irrelevant information in the description worsens the condition number of the design matrix and the numerical stability of the problem.
@@
Particularly when considering long-range descriptors, it is crucial to carefully select the information to be included. Ideally, all hyperparameters that are relevant for the exponents of the exponentially decaying radial basis functions in the density-based approach should be optimized through a grid search for each material separately.
III.3 Zirconia
Bulk zirconia ($\text{ZrO}_{\text{2}}$) is another material with a significant difference in electronegativity between its two atom types. It is used as a second test material for the descriptors developed here. The combined descriptors and the LODE method are used to assess the model’s performance.

### v3

@@ III.1 Long-Range Effects of Point Charges
As an initial test set for the present long-range model, a gas of randomly distributed point charges, analogous to the one used in Ref. Grisafi and Ceriotti, 2019, was constructed. The set comprises systems of varying volumes, each containing 64 atoms, with 32 having a positive charge of $+1$ and the other 32 having a negative charge of $-1$. To prevent large energies the minimum distance between two atoms was set to $2.5$. The training and validation data were generated using the Ewald energies and forces (as implemented in VASP). Only long-range descriptors were used for this system.
@@
Figure 1: The RMSPE in the forces for various values of the minimal $\zeta_{\text{min}}=\zeta_{\text{max}}/s^{N_{\text{max}}-1}$ of the radial basis functions, scaling constants $s$, and number of radial basis functions $N_{\text{max}}$ for the gas of point charges. It is evident that the error reduction is primarily associated with a decreasing scaling constant. All curves with a small scaling constant of $s=2$ yield a minimum at $\zeta_{\text{min}}\approx 0.09$. Exponents larger than this minimum value are insufficiently long-ranged to capture long-range electrostatic interactions accurately.
@@
Overall, reducing the scaling constant $s$ reduces the RMSPE. This is because a smaller scaling constant results in a denser distribution of radial basis functions, allowing for better a description of the electrostatic kernel ($1/r$). For the electrostatic interactions to be accurately described, it is necessary to have a sufficient number of basis functions such that they decay sufficiently slowly. We see a more or less pronounced minimum around $\zeta_{\text{min}}\approx 0.09$. The second minimum to the left corresponds to the case that the second smallest $\zeta$ becomes approximately $0.09$. For $\zeta$, the inclusion of values smaller than $0.09$ is unnecessary for this specific training set but also does not significantly degrade the quality of the fit.
The plot shows a steep increase in error on the right-hand side, indicating that the smallest exponent is insufficient to capture relevant long-range electrostatic interactions. This occurs when the radial basis function with the smallest exponent decays too quickly and the basis functions are unable to model the relevant interactions.
@@
The values $N_{\text{max}}=6$, $s=2$ and $\zeta_{\text{max}}=2$ are the optimal choices among all possible values for this specific test system, but depend somewhat on the cell size and shortest distance. These hyperparameters will be used in the subsequent calculations. Similar tests must be conducted to determine the optimal set of these three hyperparameters for other materials. However, the values obtained here provide a good starting point, which can help avoid extensive grid searches.
Next, we compare our approach with the LODE method, which is implemented in the way presented in II.4. It is known from theoretical considerations that the two-body LODE descriptor is ideal for a system where only Coulomb interactions are present Grisafi and Ceriotti (2019). Consequently, using only two-body descriptors and a single radial basis function for a small cutoff is sufficient. In this case one descriptor for each interaction pair (+1, +1), (+1, -1), (-1, +1), and (-1, -1) suffices, as these exactly represent the interaction of one positive or negative point charge with all other positive and negative point charges. In our case, the radial basis function, onto which the potential is projected, is a spherical Bessel function that is smoothly cut off at the cutoff radius.
@@
Figure 2 shows the learning curves for the two different methods. The LODE method was tested with various combinations of cutoff radii $R_{\text{cut}}$ and number of radial basis functions $N_{\text{max}}$, resulting in different numbers of expansion coefficients, and potentially probing the electrostatic potentials at points further away from the central atom. For the LODE model, this adds irrelevant information for this specific simple $1/r$ point charge model.
@@
where $\textbf{r}_{i}$ and $\textbf{r}_{j}$ are the positions of the particles and $Z_{i}$ and $Z_{j}$ are their charges. Since the descriptors include information about each atom type $J$ the charges are not relevant for the implementation and can be set to one. Then the regression will essentially determine $Z_{i}Z_{j}$.
@@
is obtained as a function of the electrostatic potential at $\textbf{r}_{i}$
@@
This potential is exactly "measured" by the LODE descriptor with one radial basis function if the cutoff radius is smaller than the minimal atomic distance (the Gaussian broadening adds additional width to the source charges though).Grisafi and Ceriotti (2019) The use of a cutoff sphere is possible due to Gauss’s law: the electrostatic potential at the center of the sphere can be determined by a probe charge of finite size, as long as there is no source term inside the sphere of the probe charge.
@@
With a cutoff radius of 3 Å, the value measured by the single descriptor does not correspond exactly to the value of the potential at the center of the sphere. Hence, the electrostatic potential of point charges that enter the projection sphere can not be determined accurately. This is why the blue line levels off at relatively high RMSPEs. Generally, considering a Gaussian broadening of $\sigma=0.3$ for the atomic source charges, a cutoff radius of $R_{\text{cut}}\lesssim 2.5~{}$Å$-5\sigma$ must be chosen, where 2.52.52.5~{}2.5Å is the shortest nearest neighbor distance in our models and the factor 5 before $\sigma$ assures that the broadened charge has decayed to negligible values.
@@
When using four radial basis functions (turquoise line), a linear combination of the values of the descriptors can approximate a $\delta$-function at the origin, although the model requires additional training data to learn the precise linear combination that corresponds best to the $\delta$-function.
@@
When training only on energies and predicting forces (as shown in Fig. 2 right panels (c) and (d)), the energy predictions are mostly more accurate. However, the errors in the predicted forces are notably higher and seem to stagnate at some point. The same precision as in the mixed training on energies and forces is never reached. This indicates that the models require information about the forces to make highly accurate predictions for forces possible.
@@
A comparison of our density-based descriptor (black line) with the ideal LODE descriptor for point charges reveals that the LODE descriptor (red line) is superior. While our density-based descriptor yields satisfactory results, it is necessary to employ more radial basis functions (specifically, six instead of one) and perform a hyper-parameter search to identify the optimal choice of $\zeta_{\text{max}}$, scaling constant $s$, and the number of radial basis functions. This shows that our density-based descriptor, as predicted by theory, is not the best possible descriptor for this prototypical system. Nevertheless, it is still capable of accurately describing the interactions, albeit requiring more training data.
@@
The LODE descriptor with one radial basis function and a cutoff radius smaller than the minimal atomic distance (red line) and the density-based descriptor (black line) can both achieve errors below 1% when sufficient training data are used. The LODE descriptor requires only about 40 training structures to achieve a relative error of 1%. The density-based descriptor requires 250 training structures to achieve the same accuracy and then stagnates. We believe the stagnation at small errors is likely related to the condition number of the design problem becoming very large, which makes it difficult for the pseudo-inversion to separate "noise" from relevant data. Nevertheless, both models give very satisfactory results. These are better than those typically obtained using machine learning, as the errors for MLFFs are typically in the mid-single digit percentage range (around 2 to 10%). The hyperparameters, such as the number of radial basis functions and reciprocal lattice points, can be adjusted to increase the speed and memory efficiency of the computations while still achieving excellent results.
III.2 Liquid Sodium Chloride
@@
After demonstrating the models’ ability to describe systems with purely $1/r$ interactions, they were applied to real materials. Liquid sodium chloride (NaCl) was chosen as the first material due to the significant difference in electronegativity between the Na and Cl atoms suggesting the presence of non-negligible long-range electrostatic interactions. The combined descriptors capture both short- and long-range interactions, enabling analysis of material properties depending on both types of interactions.
@@
The dataset for this material consists of 1014 different structures, each containing 64 Na and 64 Cl atoms. The configurations are taken from VASP molecular dynamics (MD) simulations with 50000 steps and a time step of 1.5 fs. During the MD, the material was heated from 1100K to 1400K using a Langevin thermostat. An energy cutoff of 350 eV was used. The PAW potentials used were "PAW_PBE Na_pv 19Sep2006" for Na and "PAW_PBE Cl 06Sep2006" for Cl, considering seven valence electrons for each element.
@@
Table 1 presents the results for NaCl obtained using five different methods: purely local/short-range descriptors (Local), MACE without MP (MACE no MP), LODE combined with short-range descriptors (LODE), long-range density-based descriptors combined with short-range descriptors (Density) and invariant MACE with one MP layer (MACE inv.) and equivariant MACE with one MP layer (MACE equiv.). The combination of short- and long-range descriptors follows the procedure outlined in Eq. (15).
@@
For the short-range part of the descriptor, the hyperparameters were set to $\sigma=0.5$, $R_{\text{cut}}=~{}$6 Å, $N_{\text{max}}=6$ and $L_{\text{max}}=3$ where $L_{\text{max}}$ corresponds to the angular quantum number and indicates the number of angular basis functions. For the density-based part, the optimal choice was $N_{\text{max}}=7$, $s=2$, $\zeta_{\text{max}}=1.5$ and $\sigma=0.3$.
For the LODE minimal, we used a model that can describe $1/r$ interactions exactly ($R_{\text{cut}}=$ 1 Å, $N_{\text{max}}=1$ and $\sigma=0.3$) with the long-range part being weighted 40 times more than the short-range part and the forces being weighted 3 times stronger than the energies. A more flexible LODE model was also used, with a larger cut-off radius of $R_{\text{cut}}=$7 Åand $N_{\text{max}}=5$, where the long-range part was weighted by a factor of 0.003 compared to the short-range part (the LODE descriptors show huge fluctuations in the case of a liquid).
A polynomial kernel of order $\zeta=4$ was used.
@@
For MACE $R_{\text{cut}}=~{}$6 Å, descriptors with up to four-body interactions, force weights of $1000$, energy weights of $10$ and $100$ epochs were used. After $80$ epochs the energy weights are increased to further lower the error in the energies.
@@
Table 1: The root-mean-square percentage error (RMSPE) was calculated for the validation data of liquid NaCl using combined density and radial LODE descriptors describing the electrostatic potential and a more flexible one as well as local-only descriptors and MACE Batatia et al. (2022) with and without Message Passing (MP) as well as using invariant (inv.) and equivariant (equiv.) features. Radial cutoffs of typically 6 Å are used. The results show that the density-based approach outperforms the LODE method for real materials and is as effective as MACE with invariant MP. Nevertheless, including long-range interactions is essential as the purely local scheme and MACE without MP yield large errors.
RMSPE E (%) RMSPE F (%)
Local 7.7 12.6
MACE no MP 8.5 8.5
LODE minimal 5.6 3.7
LODE flexible 7.0 11.0
Density 2.2 3.3
MACE inv. 3.0 3.1
MACE equiv. 2.1 2.2
The density-based descriptors and Message Passing MACE outperform the purely local method. LODE performs better than the local methods but not as good as the density-based and MP models.
The likely explanation for the better performance of the density-based approach is that the LODE is optimal for purely electrostatic interactions, yet lacks the capability to describe intermediate range interactions. In real materials, the $Z_{i}Z_{j}/\|\textbf{r}_{ij}\|=Z_{i}Z_{j}/r_{ij}$ Coulomb interaction between particles $i$ and $j$ is screened by the electrons, resulting in an effective interaction
@@
At very large distances the interaction will be screened by the ion-clamped macroscopic dielectric constant $\epsilon^{-1}$. However, at medium distance, the electronic screening is less pronounced and the interaction approaches $1/r$. LODE is designed to model a functional form of $1/r$ only but can, by construction, not model screened interactions that deviate from the functional $1/r$ behavior. The present approach is more flexible as the functional form of the interaction is not keyed into the model but rather learned.
Even increasing the cutoff radius of the LODE method to 7 Å and using more radial basis functions (LODE flexible) does not improve the results. On the contrary, the errors in the forces increase by a factor 1.3, indicating that the electrostatic potential at points further away from the atom of interest does not add any relevant information to the representation of the atomic environment.
@@
Even when the cutoff radius $R_{\text{cut}}$ of the purely local method is systematically increased to make the descriptors longer-ranged the result never becomes quite as good as for the combined density method or MACE. This is shown in Fig. 3. These data were calculated using the MLFF implemented in VASP. For the three-body descriptors a cutoff radius of $5$ was used, since this yields better results than 6 Å. This is also the reason why the initial value in this plot is better than the one shown in Tab. 1 obtained using polipy4vasp. In polipy4vasp a distinction between two-body and three-body cutoffs is impossible and the best compromise is found to be 6 Å. The initial error (9.8%) is comparable to MACE without MP (8%) but requires us to perform a hyperparameter search for the cutoff of the three-body terms. As shown in Fig. 3, the error decreases as the cutoff of the two-body descriptors increases, reaching a value slightly below 5 % when the cutoff is sufficiently large. In summary, naively increasing the two-body cutoff does not quite allow us to recover the accuracy of the long-range models. Nevertheless, the present observations have prompted us to routinely use relatively larger cutoffs for the two-body descriptors (8 Å), and fairly small radial cutoffs for the three-body descriptors (5 Å) in VASP as default.
@@
Figure 3: RMSPE in the forces as a function of the two-body cutoff radius $R_{\text{cut}}$ for NaCl for the short-range descriptors. Data was calculated using VASP. The local method cannot match the accuracy of the combined density method or MACE even when the cut-off radius is increased to 12 Å.
@@
Furthermore, irrelevant information worsens the condition number of the design matrix and the numerical stability of the problem.
@@
Especially when considering long-range descriptors, it is crucial to carefully select the information to be included. Ideally, all relevant hyperparameters should be optimized by a grid search.
III.3 Solid Sodium Chloride
As a less demanding test, we investigate solid NaCl. To construct a training set, a 128 atom fcc supercell was heated from 200 K to 900 K in 100,000 steps (1.5 ns) for an NPT ensemble. The same hyperparameters as for liquid sodium chloride were used to fit the data, except that the forces and energies were weighted equally.
Table 2 shows that the errors for forces and energies are about a factor of three smaller than for liquid NaCl, a finding we typically find: liquids are much more difficult to fit than solids. Again, the single-layer MACE model and the local model show rather similar errors, and the inclusion of long-range density descriptors reduces the error by almost a factor of 3. In this case, LODE and the density descriptors show remarkably similar errors. As for liquid NaCl, increasing the cut-off radius for LODE and the number of basis functions worsens the results.
The performance of MACE with invariant and equivariant message passing is even better than for the liquid, improving by a factor of 2 compared to both LODE and the long-range density descriptors. An error well below 1 % is indeed quite remarkable.
Table 2: The root-mean-square percentage error (RMSPE) was calculated for the validation data of solid NaCl using combined density and radial LODE descriptors describing the electrostatic potential and a more flexible one as well as local-only descriptors and MACE Batatia et al. (2022) with and without Message Passing (MP) as well as using invariant (inv.) and equivariant (equiv.) features. Radial cutoffs of typically 6 Å are used. The results show that LODE and the density-based approach outperform the purely local method for real materials and including long-range interactions is essential.
RMSPE E (%) RMSPE F (%)
Local 0.7 4.2
MACE no MP 0.8 3.8
LODE minimal 0.4 1.4
LODE flexible 0.6 2.6
Density 0.4 1.5
MACE inv. 0.3 0.7
MACE equiv. 0.3 0.7
III.4 Zirconia
Bulk zirconia ($\text{ZrO}_{\text{2}}$) is a material with a significant difference in electronegativity between its two atom types, but also exhibits strong covalent bonding as opposed to NaCl. It hence poses a more significant challenge than a simple ionic material.